# Deterministic Dynamic Edge-Colouring

### v1

@@ 5.3 Correctness
Claim 36.
For $10<\varepsilon<$ 1 and $0<\eta\leq\tfrac{\varepsilon}{128\log n}$, we have that $(1+\eta)^{4\log n}\leq(1+\tfrac{\varepsilon}{16})$
Proof.
It follows from the well-known fact that $1+r\leq e^{r}$ for all $r\in\mathbb{R}$ that $(1+\tfrac{x}{q})^{q}\leq e^{x}$ for all reals $x,q\neq 0$. Indeed, we can write;
$(1+\eta)^{4\log n}\leq(1+\tfrac{\varepsilon}{128\log n})^{4\log n}\leq e^{\tfrac{\varepsilon}{32}}$
Furthermore, for $r>-1$ it is well-known that $\tfrac{r}{1+r}\leq\log(1+r)$ and so we have
$(1+\eta)^{4\log n}\leq e^{\tfrac{\varepsilon}{16(1+\tfrac{\varepsilon}{16})}}\leq(1+\tfrac{\varepsilon}{16})$
as claimed. ∎
@@
A similar argument shows that we perform at most $2\sqrt{\log n}$ splits with $t=t_{2}$ before $\tfrac{2^{10\sqrt{\log n}}}{t_{2}^{i-1}}$ goes below $1/2$. By Lemma 36, we have $(1+\eta)^{4\sqrt{\log n}}<2$, and so $\hat{\Delta}_{4\sqrt{\log n}}<1$, contradicting the fact that $\hat{\Delta}_{2\sqrt{\log n}}$ must upper bound the maximum degree of all graphs in $\mathcal{G}_{4\sqrt{\log n}}$, one of which must have degree at least one, since $\mathcal{G}_{2\sqrt{\log n}}$ partitions $G$. ∎
Finally, we show that any $(G,\Delta_{\max},\varepsilon)$-valid-hierarchy with $\Delta_{\max}\geq\tfrac{10^{7}\log^{5}n}{\varepsilon^{3}}$ satisfies the conditions of Theorem 4. We note that if $\Delta_{\max}<\tfrac{10^{7}\log^{5}n}{\varepsilon^{3}}$, we do not need to apply the hierarchy, as we can simply apply the techniques develop in Section 6 directly on the graph.
Lemma 38.
Given any $(G,\Delta_{\max},\varepsilon)$-valid-hierarchy with $\Delta_{\max}\geq\tfrac{10^{7}\log^{5}n}{\varepsilon^{3}}$. Let $i$ be the smallest integer $i$ for which $\tfrac{\Delta_{\max}}{t_{1}^{i}}<\lfloor 2^{10\sqrt{\log n}}\rfloor$. Similarly, let $j$ be the smallest integer $j$ for which $\tfrac{\Delta_{\max}}{t_{1}^{i}t_{2}^{j}}<\tfrac{10^{7}\log^{5}n}{\varepsilon^{3}}$ Then, assuming the hierarchy has depth $i+j$, we have:
@@
    $\hat{\Delta}_{i+j}=(1+\eta)^{i+j}\tfrac{\Delta_{\max}}{t_{1}^{i}t_{2}^{j}}\leq(1+\tfrac{\varepsilon}{16})\tfrac{\Delta_{\max}}{t_{1}^{i}t_{2}^{j}}$
Proof.
We show the first item first. Observe that we perform $i$ splits into at most $t_{1}$ graphs first, meaning that $|\mathcal{G}_{i}|\leq t_{1}^{i}$ by an easy induction on $i$, since $|\mathcal{G}_{0}|=1$. A similar induction on $j$ yields the result, since this time each split is into $t_{2}$ graphs. The proof of the first part of the second item is a completely synchronous induction on first $i$ and then $j$ using item 5 of Definition 31.
The last in-equality of item $2$ follows by observing that $i+j\leq 4\sqrt{\log n}$ by Lemma 37 and then applying Claim 36. ∎
Lemma 35 and Lemma 38 together establish correctness of the Algorithm. In order to prove Theorem 4, we need now only analyse the running time and the recourse of the hierarchy. We do so in the next sub-section.
@@ 5.4 Analysis
Lemma 39.
Starting from an empty graph, a sequence of $\sigma$ insertion or deletions to graphs at level $i$ triggers at most $\sigma\mathopen{}\left(2^{12}\varepsilon^{-2}\log n\right)\mathclose{}^{j-i}$ insertions or deletions to graphs at level $j\geq i$.
Proof.
The proof is by strong induction on $j-i$. The base case where $j-i=0$ clearly holds. We proceed to the induction step. Assume that for $j-i=0$, $j-i=1$, …, $j-1=k$ the lemma holds. Now assume $j-i=k+1$. First of all, we note that any insertion or deletion at a lower level might trigger a similar insertion or deletion at all higher levels. By the induction hypothesis, the number of such insertion or deletions, $U_{1}$, at level $j$ is bounded as follows:
$U_{1}\leq\sum\limits_{\ell=i}^{j-1}\sigma\mathopen{}\left(2^{12}\varepsilon^{-2}\log n\right)\mathclose{}^{\ell-i}\leq\tfrac{\sigma}{2}\mathopen{}\left(2^{12}\varepsilon^{-2}\log n\right)\mathclose{}^{j-i}$
since $\sum_{i=1}^{y}x^{i}\leq\tfrac{x^{y+1}}{x-1}$ for $x\geq 2$. Note that these insertions and deletions account not only for direct insertions or deletions into $G$, but also for insertions and deletions resulting from re-colourings at some level $\ell<j-1$. Indeed, these re-colourings trigger insertions and deletions at higher levels, whose contributions are already accounted for in the quantity above.
Thus it only remains to account for the number of insertions or deletions, $U_{2}$, caused by re-colourings at level $j-1$. Observe that by Lemma 27 it follows that over any sequence of $\sigma^{\prime}$ insertions or deletions to a graph at level $j-1$, at most $\sigma^{\prime}\cdot{}\tfrac{2^{10}\log n}{\varepsilon^{2}}$ recolourings are performed at level $j-1$. Each recolouring corresponds to one insertion and one deletion on the next level, and so we find that at most
@@
insertions or deletions are performed at level $j$ as a result of re-colourings at level $j-1$. Summing $U_{1}$ and $U_{2}$ finishes the induction. ∎
@@
Proof.
Starting from any empty graph undergoing a sequence of $\sigma$ updates Lemma 39 and Lemma 37 together imply that at most $\sigma\cdot{}2^{\tilde{O}_{\log\varepsilon^{-1}}(\sqrt{\log n})}$ insertion and deletions are performed at level $\ell$ for any $\ell\leq h$. Since the hierarchy maintained has depth $h\leq 4\sqrt{\log n}$ by Lemma 37, an insertion or deletion triggers at most
$\sigma\cdot{}2^{\tilde{O}_{\log\varepsilon^{-1}}(\sqrt{\log n})}$
re-colourings, insertion or deletions to some level of the hierarchy. Hence, over the sequence of $\sigma$ updates the recourse is bounded by $\sigma\cdot{}2^{\tilde{O}_{\log\varepsilon^{-1}}(\sqrt{\log n})}$.
@@
By Theorem 3, we can process a sequence of $\sigma_{v}$ insertions or deletion to a graph of level $|v|$ in $c\sigma_{v}\log(m)t_{1}^{3}$ update time for some constant $c$. As a result, the total update time is
@@
and so the amortised update time is as claimed. ∎
[…]
@@ 6 Deterministic Dynamic Edge-Colouring of Low-Degree Graphs
In this section, we show how to deterministically maintain a $(\Delta_{\max}+1)$-edge-colouring in a dynamic graph with maximum degree $\Delta\leq\Delta_{\max}$. In Section 7.2, we will show how to adjust the algorithm to the case where no upper bound on $\Delta$ is needed, at the cost of the algorithm using $(1+\varepsilon)\Delta$ colours instead. We will follow the approach sketched in the informal overview. We begin by showing some structural corollaries of previous work:
6.1 Structural Corollaries
As mentioned earlier, a key observation is that a particular bichromatic path cannot be reached from too many different starting points. Similar observation has been used in the literature before [Ber22, BD23, Chr23] to analyse algorithms. Here, we will apply the fact more explicitly and actually update data structures at every point that may reach a recoloured edge via a short and non-overlapping multi-step process with few steps. We will use the following Lemma due to Christiansen [Chr23]. Note that originally the lemma is stated using the maximum degree $\Delta$, but the Lemma carries over ad-verbatim if one replaces $\Delta$ by an upper bound $\Delta_{\max}$ on $\Delta$:
Lemma 41 (Lemma 26 in full-version of [Chr23]).
Let $H\subset G$ be a subgraph of $G$ and let $c,c^{\prime}$ be two (partial) $(\Delta_{\max}+1)$-edge-colourings of $G$ differing only on $H$. Then, under both $c$ and $c^{\prime}$, the same at most $2(\Delta_{\max}+1)^{3}\cdot{}|H|$ points in $G-N^{1}(H)$ can reach $H$ via a 1-step process.
We have the following Corollary 42 below:
Corollary 42.
Let $G$ be a graph with maximum degree $\Delta\leq\Delta_{\max}$ endowed with a proper (possibly partial) $(\Delta_{\max}+1)$ edge-colouring $c$. Let $v\in V(G)$ be any vertex. Then $v$ can be reached by at most $3^{i}(\Delta_{\max}+1)^{3i}$ non-overlapping $i$-step processes.
Furthermore, if $C$ is a semi-consistent $j$-step process with step-length $\leq\ell$, then $N^{2}(C)$ can be reached by at most $3^{i}(\Delta_{\max}+1)^{3i+2}\cdot{}j\cdot{}\ell$ non-overlapping $i$-step processes.
Proof.
We show both claims via induction on $i$. We begin with the first claim. To this end, fix an arbitrary vertex $v$. Then $v$ can be reached by at most $2(\Delta_{\max}+1)^{3}$ points in $G-N^{1}(v)$ via a bichromatic path by Lemma 41. Any point in $N^{1}(v)$ might also reach $v$ by simply hopping to a neighbour, and so we conclude that in total at most $2(\Delta_{\max}+1)^{3}+(\Delta_{\max}+1)\leq 3(\Delta_{\max}+1)^{3}$ points are able to reach $v$ by a 1-step process. This concludes the base case. Now let $w$ be some point that can reach $v$ via $(i-1)$-steps of the above process. Then by applying the same argument as above, we find that at most $3(\Delta_{\max}+1)^{3}$ points can be the basis of a $1$-step process that can reach $w$. By applying induction we find that there are at most $3(\Delta_{\max}+1)^{3(i-1)}$ choices for $w$, and so since any $i$-step process reaching $v$ has to go through $w$, we find that there are at most
$3(\Delta_{\max}+1)^{3}\cdot{}3^{i-1}(\Delta_{\max}+1)^{3(i-1)}=3^{i}(\Delta_{\max}+1)^{3i}$
such points, which is what we wanted to show.
The second claim follows a similar argument to above. We will let $H=N^{2}(C)$ be the one-hop neighbourhood of the $j$-step process, and focus on the first vertex in $H$ reached by some process. Then we can apply Lemma 41 and the same argument as above to find that at most $3(\Delta_{\max}+1)^{3}|H|$ points can reach $H$ via a $1$-step process. The induction step is similar to above, except for the fact that we now have at most $3^{i-1}(\Delta_{\max}+1)^{3(i-1)}|H|$ choices for $w$, and so this time we find that there at most
$3(\Delta_{\max}+1)^{3}\cdot{}3^{i-1}(\Delta_{\max}+1)^{3(i-1)}|H|=3^{i}(\Delta_{\max}+1)^{3i}|H|=3^{i}(\Delta_{\max}+1)^{3i+2}\cdot{}j\cdot{}\ell$
of the desired points, which is what we wanted to show. ∎
Next we will need the following notation. Note that the definitions deviate slightly from those presented in [Chr23], as we now require the Vizing chains to be consistent according to Definition 14.
Definition 43 ([Chr23]).
Given a proper partial colouring and an uncoloured edge $e$, let $v\in R_{i}(e,\ell)$ denote the set of vertices such that there exists a consistent and non-overlapping $i$-step Vizing chain from $e$ ending at $v$ such that every Vizing chain in the $i$-step Vizing chain has length at most $\ell$.
Definition 44 ([Chr23]).
Given a proper partial colouring and an uncoloured edge $e$, let $v\in N_{i}(e,\ell)\subset R_{i}(e,\ell)$ if there exists a consistent and non-overlapping $i$-step chain from $e$ ending at $v$, and, furthermore, the multi-step Vizing chain may be extended to a non-overlapping $(i+1)$-Vizing chain by concatenating a Vizing chain of length at most $\ell$ at $v$.
In order to make it easier to maintain the data structures used to find short augmenting multi-step Vizing chains, we have imposed some further restrictions on how the chains are allowed to be extended. Most of these restrictions are already implicitly applied in for instance Christiansen’s analysis of multi-step Vizing chains [Chr23]. While this analysis allow for Vizing chains not adhering to the restrictions, we have imposed, all arguments essentially impose most of these restrictions on the chains anyway, and so the arguments go through with the above restrictions added almost ad verbatim. Hence, the following lemma is implicit in [Chr23]:
Lemma 45 (Implicit in Lemma 27 in full-version of [Chr23]).
Suppose we are given a proper partial colouring and an uncoloured edge $e$ in a graph of maximum degree $\Delta\leq\Delta_{\max}$, and let $\ell\geq 80(\Delta_{\max}+1)^{7}$. Suppose that for all $11\leq j\leq i-$ 1 it holds that no vertex in $N_{j}(e,\ell)$ can be extended to an augmenting and non-overlapping $(j+1)$-step Vizing chain via a Vizing chain of length at most $\ell$. Suppose furthermore that the conditions below hold for all $11\leq j\leq i-$ 1:

### v2

@@ 5.3 Correctness
Claim 5.7.
For $0<\varepsilon<1$ and $0<\mu\leq\tfrac{\varepsilon}{128\log n}$, we have that $(1+\mu)^{4\log n}\leq(1+\tfrac{\varepsilon}{16})$
Proof 14.
It follows from the well-known facts that $1+r\leq e^{r}$ for all $r\in\mathbb{R}$ and that $(1+\tfrac{x}{q})^{q}\leq e^{x}$ for all reals $x,q\neq 0$. Indeed, we can write;
$(1+\mu)^{4\log n}\leq(1+\tfrac{\varepsilon}{128\log n})^{4\log n}\leq e^{\tfrac{\varepsilon}{32}}$
Furthermore, for r>1r>-1 it is well-known that $\tfrac{r}{1+r}\leq\log(1+r)$ and so we have
$(1+\mu)^{4\log n}\leq e^{\tfrac{\varepsilon}{16(1+\tfrac{\varepsilon}{16})}}\leq(1+\tfrac{\varepsilon}{16})$
as claimed.
@@
A similar argument shows that we perform at most $2\sqrt{\log n}$ splits with $t=t_{2}$ before $\tfrac{2^{10\sqrt{\log n}}}{t_{2}^{i-1}}$ goes below $1/2$. By Lemma 5.7, we have $(1+\mu)^{4\sqrt{\log n}}<2$, and so $\hat{\Delta}_{4\sqrt{\log n}}<1$, contradicting the fact that $\hat{\Delta}_{4\sqrt{\log n}}$ must upper bound the maximum degree of all graphs in $\mathcal{G}_{4\sqrt{\log n}}$, one of which must have degree at least one, since $\mathcal{G}_{2\sqrt{\log n}}$ partitions $G$.
Finally, we show that any $(G,\Delta_{\max},\varepsilon)$-valid-hierarchy with $\Delta_{\max}\geq\tfrac{10^{7}\log^{5}n}{\varepsilon^{3}}$ satisfies the conditions of Theorem 1.4. We note that if $\Delta_{\max}<\tfrac{10^{7}\log^{5}n}{\varepsilon^{3}}$, we do not need to apply the hierarchy, as we can simply apply the techniques developed in Section 6 directly on the graph.
Lemma 5.9.
Given any $(G,\Delta_{\max},\varepsilon)$-valid-hierarchy with $\Delta_{\max}\geq\tfrac{10^{7}\log^{5}n}{\varepsilon^{3}}$. Let $i$ be the smallest integer $i$ for which $\tfrac{\Delta_{\max}}{t_{1}^{i}}<\lfloor 2^{10\sqrt{\log n}}\rfloor$. Similarly, let $j$ be the smallest integer $j$ for which $\tfrac{\Delta_{\max}}{t_{1}^{i}t_{2}^{j}}<\tfrac{10^{7}\log^{5}n}{\varepsilon^{3}}$. Then, if the hierarchy has depth $i+j$, we have:
@@
    $\hat{\Delta}_{i+j}=(1+\mu)^{i+j}\tfrac{\Delta_{\max}}{t_{1}^{i}t_{2}^{j}}\leq(1+\tfrac{\varepsilon}{16})\tfrac{\Delta_{\max}}{t_{1}^{i}t_{2}^{j}}$
Proof 16.
We show the first item first. Observe that we perform $i$ splits into at most $t_{1}$ graphs first, meaning that $|\mathcal{G}_{i}|\leq t_{1}^{i}$ by an easy induction on $i$, since $|\mathcal{G}_{0}|=1$. A similar induction on $j$ yields the result, since this time each split is into $t_{2}$ graphs. The proof of the first part of the second item is a completely synchronous induction on first $i$ and then $j$ using item 5 of Definition 5.2.
The last in-equality of item $2$ follows by observing that $i+j\leq 4\sqrt{\log n}$ by Lemma 5.8 and then applying Claim 5.7.
Lemma 5.6 and Lemma 5.9 together establish correctness of the Algorithm. In order to prove Theorem 1.4, we need now only analyse the running time and the recourse of the hierarchy. We do so in the next sub-section.
@@ 5.4 Analysis
Lemma 5.10.
Starting from an empty graph, a sequence of $\sigma$ insertion and deletions to graphs at level $i$ triggers at most $\sigma\mathopen{}\left(2^{12}\varepsilon^{-2}\log n\right)\mathclose{}^{j-i}$ insertions and deletions to graphs at level $j\geq i$.
Proof 17.
The proof is by strong induction on jij-i. The base case where ji=0j-i=0 clearly holds. We proceed to the induction step. Assume that for ji=0j-i=0, ji=1j-i=1, …, ji=kj-i=k the lemma holds. Now assume ji=k+1j-i=k+1. A deletion or an insertion can happen in layer $j$ only if an insertion or deletion is performed at this or an earlier layer, or if a recolouring is performed at an earlier layer.
By the induction hypothesis, the number of insertions and deletions, $U_{1}$, at level $j$ resulting from an insertion or an deletion at an earlier layer can be bounded as follows by applying the induction hypothesis:
$U_{1}\leq\sigma\mathopen{}\left(2^{12}\varepsilon^{-2}\log n\right)\mathclose{}^{j-i-1}\leq\tfrac{\sigma}{2}\mathopen{}\left(2^{12}\varepsilon^{-2}\log n\right)\mathclose{}^{j-i}$
It remains to account for the number of insertions or deletions, $U_{2}$, caused by re-colourings at level j1j-1. Note that any recolourings at level j2j-2 or higher results in deletion at level j1j-1, and are therefore accounted for in $U_{1}$. Observe that by Lemma 4.9 it follows that over any sequence of $\sigma^{\prime}$ insertions and deletions to a graph at level j1j-1, at most $\sigma^{\prime}\cdot{}\tfrac{2^{10}\log n}{\varepsilon^{2}}$ recolourings are performed at level j1j-1. Each recolouring corresponds to one insertion and one deletion on the next level, and so we find that at most
@@
insertions or deletions are performed at level $j$ as a result of re-colourings at level j1j-1. Summing $U_{1}$ and $U_{2}$ finishes the induction.
@@
Proof 18.
Starting from any empty graph undergoing a sequence of $\sigma$ updates Lemma 5.8 and Lemma 5.10 together imply that at most $\sigma\cdot{}2^{\tilde{O}_{\log\varepsilon^{-1}}(\sqrt{\log n})}$ insertion and deletions are performed at level $\ell$ for any $\ell\leq h$, since the hierarchy maintained has depth $h\leq 4\sqrt{\log n}$. An insertion or deletion triggers at most $O(\tfrac{\log n}{\varepsilon^{2}})$ re-colourings, so at most $\sigma\cdot{}2^{\tilde{O}_{\log\varepsilon^{-1}}(\sqrt{\log n})}$ recourse happens at each level. Hence, over the sequence of $\sigma$ updates the total recourse is bounded by $\sigma\cdot{}2^{\tilde{O}_{\log\varepsilon^{-1}}(\sqrt{\log n})}$.
@@
By Theorem 1.3, we can process a sequence of $\sigma_{v}$ insertions and deletions to a graph of level $|v|$ in $c\sigma_{v}\log(m)t_{1}^{3}$ update time for some constant $c$. As a result, the total update time is
@@
and so the amortised update time is as claimed.
Finally, we note that Lemma 5.8 shows that the while-loop of $\texttt{Initialize-Parameters}(\Delta_{\max},\varepsilon)$ has few iterations. Secondly, each iteration of the while-loop can be implemented efficiently provided that all values are properly discretised as noted in Remark 5.5. The time spent on this step is therefore only $O(\sqrt{\log n})$. This step only has to be performed once upon initialization. Combining all of the above now yields Theorem 1.4.
@@ 6 Deterministic Dynamic Edge-Colouring of Low-Degree Graphs
In this section, we show how to deterministically maintain a $(\Delta_{\max}+1)$-edge-colouring in a dynamic graph with maximum degree $\Delta\leq\Delta_{\max}$. In Section 7.2, we will show how to adjust the algorithm to the case where no upper bound on $\Delta$ is needed, at the cost of the algorithm using $(1+\varepsilon)\Delta$ colours instead. We will follow the approach sketched in the informal overview. We begin by showing some theoretical and structural results, which form the basis of the algorithm:
6.1 Structural results
In this section, we will show some theoretical results about stepping processes, which will form the basis for the algorithm used for edge-colouring graphs with small maximum degree. We will fix some graph $G$ and a proper (partial) colouring $c$ of the edges of $G$. We assume that $G$ has maximum degree at most $\Delta_{\max}$, and hence we allow the case where $\Delta(G)<\Delta_{\max}$. It will be assumed in all statements that this is the case, and we will not specify it further in this section. We will also be working with parameters $\ell$, $a$, and $\gamma$. These parameters often need to satisfy different requirements, which will be specified in the required lemmas. Towards the end of the section, we will give explicit values for these parameters which will be sufficient for the needed lemmas. We begin with some basic definitions which we will use to make precise statements.
6.1.1 Representing strings and sets
Given a stepping process $C=w_{1}+y_{1}+P_{1}+\dots+w_{i}+y_{i}+P_{i}$, we can represent it as a string $s(C)=P_{1}P_{2}\dots P_{i}$ and a starting node $w(C)=w_{1}$. Here, we encode each $P_{j}$ as its own unique character in the alphabet over which the string is constructed. As such if $P$ and $P^{\prime}$ differ in at least one way, the character they represent in the string are different. Even if $P$ and $P^{\prime}$ represent the same maximal bichromatic path, but with different starting points, they will be considered different. This means that the length of the string $s(C)$ is $i$ if $C$ takes $i$ steps. We emphasize again that the order in which $P_{j}$ is listed matters, so that $y_{j}$ coincides with the first vertex of $P_{j}$. Furthermore, we apologize for the ambiguity of the notation $P_{j}$ – in strings it is viewed as a single character, outside of strings as an ordered list of vertices forming a path.
Given a set $\mathcal{S}$ of strings representing stepping processes, we let
$\mathcal{S}(p):=\{s|s\in\mathcal{S},\text{lcp}(p,s)=p\}$
i.e. we include only the strings of $\mathcal{S}$ which have $p$ as a prefix. We let
$\mathcal{S}_{i,j}:=\{s|s=\bar{s}[i\dots j]\text{ for some }\bar{s}\in S\}$
i.e. the set containing every string from $\mathcal{S}$ restricted to its substring from the $i^{\text{th}}$ to the $j^{\text{th}}$ character.
For two strings $s$ and $\bar{s}$, we say that they agree until $i$ if the length of their longest common prefix is $i$ and that they differ at $i$ if the length of their longest common prefix is exactly i1i-1.
We are interested in sets of strings which do not share prefixes, as they encode stepping process that are somewhat ’spread out’. In particular, we care about the following property:
Definition 6.1.
Given a set $\mathcal{S}$ of strings representing stepping processes all with the same starting node $w$, we say that $\mathcal{S}$ is $(a,i)$-good if the compressed trie $T$, over all strings in $\mathcal{S}$, satisfies that all nodes of depth $\leq i-1$ is either a leaf or an internal node with at least $a$ children i.e. degree at least $a+1$ (except for the root which could have degree $a$).
We will work with graded sets of the form $A=\bigcup\limits_{k=0}^{\infty}A_{k}$. For an element $\alpha\in A$, it must be the case that there exists an $i$ for which $\alpha\in A_{i}$, but it need not be the case for more than one such index. Given two graded sets $A,B$, we let $A\cup B$ be the graded set defined as $(A\cup B)_{i}=A_{i}\cup B_{i}$.
Typically, we want the sets of strings that we are working with to satisfy some additional regularity requirements. These are outlined in the definition below:
Definition 6.2.
Given a set of $\mathcal{S}$ of strings representing stepping processes all with the same starting node $w$, we say that $\mathcal{S}$ is $(i,\ell,A,B,P,y)$-restricted for neighbour $y$ and maximal bichromatic path $P$ beginning at $y$, if for all $s\in\mathcal{S}$, we have:
If $s$ represents the stepping process $w_{1}+y_{1}+P_{1}+\dots+w_{j}+y_{j}+P_{j}$, then: